### v1

@@ 10 Related Work
CTMCs. CTMCs Campbell et al. (2022) have been used for various applications in generative modeling (”discrete diffusion” models), including text and image generation Shi et al. (2024); Gat et al. (2025); Shaul et al. (2024); Sahoo et al. (2024) and molecular design Gruver et al. (2023); Campbell et al. (2024); Lisanza et al. (2024). While here we use a RND for CTMCs running in reverse time, one recovers the loss functions of these generative models considering a RND of two forward time CTMCs (in Appendix A, we show this in detail).
Transport and sampling. Over the past decade there has been continued interest in combining the statistical guarantees of MCMC and IS with learning transport maps. A non-parametric version of this is described in Marzouk et al. (2016), and a parametric version through coupling-based normalizing flows was used to study systems in molecular dynamics and statistical physics Noé et al. (2019); Albergo et al. (2019); Gabrié et al. (2022); Wang et al. (2022). These methods were extended to weave normalizing flows with SMC moves Arbel et al. (2021); Matthews et al. (2022). More recent research focuses on replacing the generative model with a continuous flow or diffusion Zhang & Chen (2022); Vargas et al. (2023); Akhound-Sadegh et al. (2024); Sun et al. (2024). Our method is inspired by approaches combining measure transport with MCMC schemes Albergo & Vanden-Eijnden (2024); Vargas et al. (2024) and other samplers relying on PINN-based objectives in continuous spaces (Máté & Fleuret, 2023; Tian et al., 2024; Sun et al., 2024).
Discrete Neural samplers. The primary alternative to what we propose is to correct using importance weights arising from the estimate of the probability density computed using an autoregressive model Nicoli et al. (2020). However, the computational cost of producing samples in this case scales naively as $O(d)$, whereas we have no such constraint a priori in our case so long as the error in the Euler sampling scheme is kept small. Other work focuses on discrete formulations of normalizing flows, but the performant version reduces to an autoregressive model Tran et al. (2019). Recent work has considered using CTMCs for sampling by parameterizing their evolution operators directly via tensor networks (Causer et al., 2025) as opposed to neural network representations of rate matrices here.
@@ 11 Experiments
As a demonstration of the validity of LEAPS in high dimensions, we sample the Gibbs distribution associated to a 2-dimensional Ising model. We choose the Ising model because it is a well-studied model. In particular, it is a solvable model, which allows us to construct a robust ground truth against which we can assess the various neural architectures underlying our algorithm. Configurations of the $L\times L$ Ising lattice follow the target distribution $\rho_{1}(x)=e^{-\beta H(x)+F_{1}}$ where $\beta$ is the inverse temperature of the system, $F_{1}$ the free energy, and $H(x):\{-1,1\}^{L\times L}\rightarrow\mathbb{R}$ is the Hamiltonian for the model defined as
$H(x)=-J\sum_{\langle i,j\rangle}x_{i}x_{j}+\mu\sum_{i}x_{i}.$ (21)
Here, $J$ is the interaction strength, $\langle i,j\rangle$ denotes summation over nearest neighbors of spins $x_{i},x_{j}$ and $\mu$ is the magnetic moment. Neighboring spins are uncorrelated at high temperature but reach a critical correlation when the temperature drops behold a certain threshold. We use LEAPS to reproduce the statistics of the theory on a $15\times 15$ lattice at parameters $\beta=0.7,J=0.4$, which approach this threshold, and compare our results against a ground truth of long-run Glauber dynamics, an efficient algorithm for simulation in this parameter regime. Note that this corresponds to a $d=15\times 15=225$ dimensional space. To make $\rho_{t}$ time dependent, we make the parameters of $J_{t}$, $\mu_{t}$, $\beta_{t}$ linear functions of time, starting from the non-interacting case $J_{0}=0$.
Results. We compare three different realizations of our method, one using LEA, and the other two using deep LEC that vary in depth. For all generated samples, we use 100 steps of integration of Equation 6. We benchmark them on the effective sample size (ESS), which is a standard measure of how many effective samples our model gives according to the variance of the importance weights (see details Appendix G). In addition, we compute various observables using the Ising configurations generated by our model, such as histograms of the magnetization $M(x)=\sum_{i}x_{i}$ compared to ground truth, as well as the two point connected correlation function
The latter is a measure of the dependency between spin values a distance $r$ in lattice separation. In Figure 4, we show in the leftmost panel that the convolutional architecture outperforms the attention-based method, and the performance gap grows as we make the LEC network deeper. In the center panel, the difference in histograms of the magnetization of lattice configurations for our models as compared to ground truth samples is shown to be statistically zero, whereas relying on local MCMC alone for the same number of sampling steps (plotted in purple) illustrates that the dynamics have not converged. In the right plot, we see clear agreement between our learned sampler and the ground truth for the connected two point function. These results show that LEAPS can be an efficient simulator of complex, high dimensional target distributions.
@@ 12 Discussion
In this work, we present the LEAPS algorithm that allows to learn a non-equilibrium sampler via CTMCs parameterized by a locally equivariant neural network that is trained to minimize the variance of proactivate IS weights. A natural direction of future work will be to connect the ideas presented here with guidance or reward fine-tuning of generative CTMC models (discrete diffusion) - a problem known to be strongly tied to sampling. Further, LEAPS could easily be extended to sample across a whole family of distributions as opposed to only for a single, fixed target. Finally, we anticipate that the use of locally equivariant neural networks as well as the proactive IS scheme presented here might be useful more broadly for probabilistic models.
13 Impact Statement
@@ 14 Acknowledgements
is a constant in time $s$. However, this constant may depend on $\mathbf{X}$. Further, note that
$\mathbb{Q}_{s}(X_{s})=|\mathcal{K}_{s}^{\theta}\rho_{s}(X_{s})-\partial_{s}F_{s}|^{2}$
@@
$\Rightarrow\quad\left[Q_{t}(x,y)-Q_{t}(y,x)\frac{q_{t}(x)}{q_{t}(y)}\right]_{+}=$ 00 0
$\Rightarrow\bar{Q}_{t}(x,y)=$ 00 0
@@
By construction, it holds that $F_{t}(\tau,i|x)$ is locally equivariant and that $[F_{t}(\tau,i|x)]_{+}=\bar{Q}_{t}(y,x)$. This finishes the proof.
@@ Appendix G Numerical experiments
Effective Sample Size
We use the self-normalized definition of the effective sample size such that, given the log weights $A_{t}$ associated to $N$ CTMC instances, the ESS at time $t$ in the generation is given by:
@@
G.1 Ising model experiments
Here we provide details of the numerical implementation of our study of the $L=15$ Ising model. For the locally equivariant attention (LEA) mechanism, we use 40 attention heads, each with query, key, and value matrices of dimension 40x40. As such, there are about 350,000 parameters in the model. In addition, the locally equivariant convolutional net (LEC) of depth three uses kernel sizes of [5, 7, 15], while the depth five version uses [3, 5, 7, 9, 15], amounting to around 100,000 parameters.

### v2

@@ 10 Related Work
CTMCs. CTMCs (Campbell et al., 2022) have been used for various applications in generative modeling (”discrete diffusion” models), including text and image generation (Shi et al., 2024; Gat et al., 2025; Shaul et al., 2024; Sahoo et al., 2024) and molecular design (Gruver et al., 2023; Campbell et al., 2024; Lisanza et al., 2024). While here we use a RND for CTMCs running in reverse time, one recovers the loss functions of these generative models considering a RND of two forward time CTMCs (see Appendix A).
Transport and sampling. Over the past decade there has been continued interest in combining MCMC and IS with learning transport maps. A non-parametric version of this is described in (Marzouk et al., 2016), and a parametric version through coupling-based normalizing flows was used to study systems in molecular dynamics and statistical physics (Noé et al., 2019; Albergo et al., 2019; Gabrié et al., 2022; Wang et al., 2022). These methods were extended to weave normalizing flows with SMC moves (Arbel et al., 2021; Matthews et al., 2022). More recent research focuses on replacing the generative model with a continuous flow or diffusion (Zhang & Chen, 2022; Vargas et al., 2023; Akhound-Sadegh et al., 2024; Sun et al., 2024). Our method is inspired by approaches combining measure transport with MCMC schemes (Albergo & Vanden-Eijnden, 2024; Vargas et al., 2024) and other samplers relying on PINN-based objectives in continuous spaces (Máté & Fleuret, 2023; Tian et al., 2024; Sun et al., 2024).
Discrete Neural samplers. The primary alternative to what we propose is to correct using importance weights arising from the estimate of the probability density computed using an autoregressive model (Nicoli et al., 2020; Wu et al., 2019; McNaughton et al., 2020). However, the computational cost of producing samples in this case scales naively as $O(d)$, whereas we have no such constraint a priori in our case so long as the error in the Euler sampling scheme is kept small. Other work focuses on discrete formulations of normalizing flows, but the performant version reduces to an autoregressive model (Tran et al., 2019). Recent work has considered using CTMCs for sampling by parameterizing their evolution operators directly via tensor networks (Causer et al., 2025) as opposed to neural network representations of rate matrices here. Finally, discrete diffusion models have already been explored in the context of unsupervised neural optimization (Sanokowski et al., 2024, 2025).
@@ 11 Experiments
As a demonstration of the validity of LEAPS in high dimensions, we benchmark it on models of statistical physics, namely the Ising model and the Potts model. We apply LEAPS on each model on a $15\times 15$ lattice corresponding to a $d=15\times 15=225$ dimensional space. To construct $\rho_{t}$, we use temperature annealing making the inverse temperature $\beta_{t}$ a linear function of time.
Ablation - Ising model. To establish the validity of our method also experimentally, we compare our results against a ground truth of long-run Glauber dynamics, an efficient algorithm for simulation in this parameter regime. In Figure 8, one can see that LEAPS recovers the expected physical observables. Further, in Figure 8 we compare three different realizations of our method, one using LEA, and the other two using deep LEC that vary in depth. The deep LEC network performs best and we use it for all subsequent experiments.
Figure 4: LEAPS vs AIS on Ising model. Comparison across different number of simulation steps for AIS. LEAPS is run for 100 steps. Note, however, that LEAPS requires neural net evaluations.
Benchmarks - Ising model. We then apply LEAPS on the parameters corresponding to the corresponding critical temperatures of the Ising model, i.e. the “hardest” temperature to sample from. Here, LEAPS achieves a high effective sample size (ESS) of $\sim 68\%$. In Figure 4, we compare LEAPS against annealed importance sampling (AIS) and see that AIS needs around $10^{6}$ integration steps to achieve a similar ESS (vs. $100$ integration steps from LEAPS). Next, we use the DISCS benchmark (Goshvadi et al., 2023) to benchmark LEAPS against MCMC samplers (see Figure 5). The ESS of LEAPS is significantly higher, even after accounting for the number of function evaluations (NFEs).
Figure 5: LEAPS vs MCMC samplers on Ising model (DISCS benchmark (Goshvadi et al., 2023)). Note that the comparison has limitations: Function call for LEAPS is neural network evaluations and energy calls for MCMC samplers. Further, ESS is measured differently for MCMC samplers.
Benchmarks - Potts model. We also apply the LEAPS method on the Potts model following Goshvadi et al. (2023). LEAPS achieves a high ESS ($\sim 20\%$) in only 100 integration steps while AIS has a lower ESS even for $10^{6}$ integration steps (see Figure 6). Also on the DISCS benchmark, LEAPS shows a higher ESS than MCMC samplers.
Overall, the above results demonstrate the validity of LEAPS and that post-training, it is able to generate realistic samples with few simulation steps, while leveraging importance weights to recover exact observables.
Figure 6: LEAPS vs AIS on Potts model for critical temperature. Comparison across different number of simulation steps for AIS. LEAPS is run for 100 steps.
@@ 12 Discussion
In this work, we present the LEAPS algorithm to learn to sample from discrete distributions via continuous-time Markov chains (“discrete diffusion”). Crucially, we translated the sampling problem into a neural network design problem. While we presented first locally equivariant neural network designs, we anticipate that future work will focus on building more scalable and more expressive locally equivariant network architectures. Another direction of future work will be to connect the ideas presented here with guidance and reward fine-tuning of generative model - a problem strongly tied to sampling. Further, neural samplers like LEAPS can be used to sample across a whole family of distributions as opposed to only for a single target.
Impact Statement
@@ Acknowledgements
is a constant in time $s$. However, this constant may depend on $\mathbf{X}$.
@@
$\Rightarrow\quad\left[Q_{t}(x,y)-Q_{t}(y,x)\frac{q_{t}(x)}{q_{t}(y)}\right]_{+}=$ $0$
$\Rightarrow\bar{Q}_{t}(x,y)=$ $0$
@@
By construction, it holds that $F_{t}(\tau,i|x)$ is locally equivariant and that $F_{t}(\tau,i|x)]_{+}=\bar{Q}_{t}(y,x)$. This finishes the proof.
@@ Appendix E Local equivariance of ConvNet
E.1 Universal representation of decomposition into locally-invariant prediction head
In this section, we answer the question: To what extend do we reduce the expressive power of the neural network by imposing the above conditions? To show a universal representation theorem, define the polynomial basis features
$P(x)=(1,x^{1},x^{2},x^{3},x^{4},\cdots)$
i.e. we assume infinite width of the neural network output here.
Proposition E.1.
Any continuous function $\tilde{F}:C\to\mathbb{R}^{d}$ defined on a compact set $C\subset\mathbb{R}^{d}$ be approximated as follows: For any $\epsilon>0$, there is a function $F=(F_{1},\cdots,F_{d}):C\to\mathbb{R}^{D}$ that can be decomposed in
$F_{i}(x)=P(x)^{T}H_{i}(x)$
for a locally-invariant function $H:C\to(\mathbb{R}^{k})^{d}$ (i.e. $H_{i}(x)$ is independent of $x_{i}$) for some $k$ such that
$\sup\limits_{x\in C}\|\tilde{F}(x)-F(x)\|<\epsilon$
In other words, for infinite width $k$, this representation is universal.
Proof.
Further, let $\tilde{F}:C\to\mathbb{R}^{d}$ be a continuous function defined on a compact set $C$. Further, write $\tilde{F}=(\tilde{F}_{1},\cdots,\tilde{F}_{d})$ for the coordinate functions. Then fix a $j$. By the Stone-Weierstrass theorem, there is a polynomial $p^{j}(x_{1},\dots,x_{d})$ of the form
$p^{j}(x_{1},\dots,x_{d})=\sum\limits_{k_{1},\dots,k_{d}\geq 0}a_{k_{1},\dots,k_{d}}^{j}x_{1}^{k_{1}}\dots x_{d}^{k_{d}}$
where $a_{k_{1},\dots,k_{d}}\neq 0$ only for a finite number of coefficients such that
$\sup\limits_{x\in C}|\tilde{F}_{j}(x)-p^{j}(x)|<\frac{\epsilon}{d}$
Then we can re-express that polynomial via
$p^{j}(x_{1},\dots,x_{d})=$ $\sum\limits_{k_{i}\geq 0}x_{i}^{k_{i}}\left(\sum\limits_{k_{i}\geq 0,i\neq j}a_{k_{1},\dots,k_{d}}\prod\limits_{j\neq i}x^{k_{j}}_{j}\right)$
$=$ $\sum\limits_{k_{i}\geq 0}P_{k_{i}}(x_{i})H_{i,k_{i}}(x)$
$=$ $P(x_{i})^{T}H_{i}(x)$
$=:$ $F_{i}(x)$
Note that $H=(H_{1},\cdots,H_{d})$ is locally invariant. Therefore, we can conclude that $F=(F_{1},\cdots,F_{d})$ fulfills the desired conditions:
$\sup\limits_{x\in C}\|\tilde{F}(x)-F(x)\|\leq\sum\limits_{i=1}^{d}\sup\limits_{x\in C}|\tilde{F}_{i}(x)-F_{i}(x)|<\sum\limits_{i}\frac{\epsilon}{d}=\epsilon$
As $\epsilon>0$ was arbitrary, this finishes the proof. ∎
@@ Appendix G Numerical experiments
All code for experiments can be found under the following link: https://github.com/malbergo/leaps/.
We briefly explain the Ising model here. The Potts model is similar and we refer to details for the Potts model to the DISCS benchmark (Goshvadi et al., 2023). Configurations of the $L\times L$ Ising lattice follow the target distribution $\rho_{1}(x)=e^{-\beta H(x)+F_{1}}$ where $\beta$ is the inverse temperature of the system, $F_{1}$ the free energy, and $H(x):\{-1,1\}^{L\times L}\rightarrow\mathbb{R}$ is the Hamiltonian for the model defined as
$H(x)=-J\sum_{\langle i,j\rangle}x_{i}x_{j}+\mu\sum_{i}x_{i}.$ (21)
Here, $J$ is the interaction strength, $\langle i,j\rangle$ denotes summation over nearest neighbors of spins $x_{i},x_{j}$ and $\mu$ is the magnetic moment. Neighboring spins are uncorrelated at high temperature but reach a critical correlation when the temperature drops below a certain threshold. As observables, we use the histograms of the magnetization $M(x)=\sum_{i}x_{i}$ compared to ground truth, as well as the two point connected correlation function
The latter is a measure of the dependency between spin values at a distance $r$ in lattice separation.
For evaluation, we use the self-normalized definition of the effective sample size such that, given the log weights $A_{t}$ associated to $N$ CTMC instances, the ESS at time $t$ in the generation is given by: